# Composing Global Optimizers to Reasoning Tasks via Algebraic Objects in Neural Nets

### v2

@@ Appendix A Decoupling $L_{2}$ Loss (Proof)
$\mathbb{E}_{y_{1},y_{2}}\left[z_{p_{1}k_{1}j_{1}}z_{p_{2}k_{2}j_{1}}z_{p_{3}k_{3}j_{2}}z_{p_{4}k_{4}j_{2}}c_{k^{\prime}j_{1}}\bar{c}_{k^{\prime}j_{2}}\phi_{k_{1}}(y_{p_{1}})\phi_{k_{2}}(y_{p_{2}})\phi_{k_{3}}(y_{p_{3}})\phi_{k_{4}}(y_{p_{3}})\right]$ (22)
@@
Case 1: $(1,3)$, $(2,4)$, $(1,4)$, $(2,3)$. The 4 cases are identical so we only need to analyze one. We take $(1,3)$ as an example. For $(1,3)$, $p_{1}=p_{3}=a$, $p_{2}=p_{4}=b$ and the only nonzero terms is when $k_{1}+k_{3}=0\mod d$, $k_{2}+k_{4}=0\mod d$, since $\mathbb{E}_{y_{1}}\left[\phi_{k_{1}}(y_{1})\phi_{k_{3}}(y_{1})\right]=\mathbb{I}(k_{1}+k_{3}=0\mod d)$ (and similar in other cases). Then Eqn. 22 becomes:
@@
$\sum_{k^{\prime}\neq 0}\sum_{j_{1}j_{2}}\sum_{p\in\{1,2\}}\sum_{k_{1}+k_{2}+k_{3}+k_{4}=0}z_{pk_{1}j_{1}}z_{pk_{2}j_{1}}z_{pk_{3}j_{2}}z_{pk_{4}j_{2}}c_{k^{\prime}j_{1}}\bar{c}_{k^{\prime}j_{2}}$ (32)
$=$ $\sum_{k^{\prime}\neq 0}\sum_{j_{1}j_{2}}\sum_{p\in\{1,2\}}\sum_{k_{1}+k_{2}=k_{3}+k_{4}}z_{pk_{1}j_{1}}z_{pk_{2}j_{1}}\bar{z}_{pk_{3}j_{2}}\bar{z}_{pk_{4}j_{2}}c_{k^{\prime}j_{1}}\bar{c}_{k^{\prime}j_{2}}$ (33)
$=$ $\sum_{k^{\prime}\neq 0}\sum_{m}\sum_{p\in\{1,2\}}\sum_{j_{1}j_{2}}\sum_{p\in\{1,2\}}\sum_{k_{1}+k_{2}=m}\sum_{k_{3}+k_{4}=m}z_{pk_{1}j_{1}}z_{pk_{2}j_{1}}\bar{z}_{pk_{3}j_{2}}\bar{z}_{pk_{4}j_{2}}c_{k^{\prime}j_{1}}\bar{c}_{k^{\prime}j_{2}}$
$=$ $\sum_{k^{\prime}\neq 0}\sum_{m}\sum_{p\in\{1,2\}}\left[\sum_{j_{1}}\left(\sum_{k_{1}+k_{2}=m}z_{pk_{1}j_{1}}z_{pk_{2}j_{1}}\right)c_{k^{\prime}j_{1}}\right]\left[\sum_{j_{2}}\left(\sum_{k_{3}+k_{4}=m}\overline{z_{pk_{3}j_{2}}z_{pk_{4}j_{2}}}\right)\bar{c}_{k^{\prime}j_{2}}\right]$

### v3

@@ Appendix A Decoupling $L_{2}$ Loss (Proof)
$\mathbb{E}_{g_{1},g_{2}}\left[z_{p_{1}k_{1}j_{1}}z_{p_{2}k_{2}j_{1}}z_{p_{3}k_{3}j_{2}}z_{p_{4}k_{4}j_{2}}c_{k^{\prime}j_{1}}\bar{c}_{k^{\prime}j_{2}}\phi_{k_{1}}(g_{p_{1}})\phi_{k_{2}}(g_{p_{2}})\phi_{k_{3}}(g_{p_{3}})\phi_{k_{4}}(g_{p_{3}})\right]$ (22)
@@
Case 1: $(1,3)$, $(2,4)$, $(1,4)$, $(2,3)$. The 4 cases are identical so we only need to analyze one. We take $(1,3)$ as an example. For $(1,3)$, $p_{1}=p_{3}=a$, $p_{2}=p_{4}=b$ and the only nonzero terms is when $k_{1}+k_{3}=0\mod d$, $k_{2}+k_{4}=0\mod d$, since $\mathbb{E}_{g_{1}}\left[\phi_{k_{1}}(g_{1})\phi_{k_{3}}(g_{1})\right]=\mathbb{I}(k_{1}+k_{3}=0\mod d)$ (and similar in other cases). Then Eqn. 22 becomes:
@@
$\sum_{k^{\prime}\neq 0}\sum_{j_{1}j_{2}}\sum_{p\in\{a,b\}}\sum_{k_{1}+k_{2}+k_{3}+k_{4}=0}z_{pk_{1}j_{1}}z_{pk_{2}j_{1}}z_{pk_{3}j_{2}}z_{pk_{4}j_{2}}c_{k^{\prime}j_{1}}\bar{c}_{k^{\prime}j_{2}}$ (32)
$=$ $\sum_{k^{\prime}\neq 0}\sum_{j_{1}j_{2}}\sum_{p\in\{a,b\}}\sum_{k_{1}+k_{2}=k_{3}+k_{4}}z_{pk_{1}j_{1}}z_{pk_{2}j_{1}}\bar{z}_{pk_{3}j_{2}}\bar{z}_{pk_{4}j_{2}}c_{k^{\prime}j_{1}}\bar{c}_{k^{\prime}j_{2}}$ (33)
$=$ $\sum_{k^{\prime}\neq 0}\sum_{m}\sum_{p\in\{a,b\}}\sum_{j_{1}j_{2}}\sum_{k_{1}+k_{2}=m}\sum_{k_{3}+k_{4}=m}z_{pk_{1}j_{1}}z_{pk_{2}j_{1}}\bar{z}_{pk_{3}j_{2}}\bar{z}_{pk_{4}j_{2}}c_{k^{\prime}j_{1}}\bar{c}_{k^{\prime}j_{2}}$
$=$ $\sum_{k^{\prime}\neq 0}\sum_{m}\sum_{p\in\{a,b\}}\left[\sum_{j_{1}}\left(\sum_{k_{1}+k_{2}=m}z_{pk_{1}j_{1}}z_{pk_{2}j_{1}}\right)c_{k^{\prime}j_{1}}\right]\left[\sum_{j_{2}}\left(\sum_{k_{3}+k_{4}=m}\overline{z_{pk_{3}j_{2}}z_{pk_{4}j_{2}}}\right)\bar{c}_{k^{\prime}j_{2}}\right]$
@@ C.3 Canonical Forms
Lemma 5 (Necessary Condition for $R_{\mathrm{c}}$).
All order-1 and order-2 solutions satisfying $R_{\mathrm{c}}:=\{r_{k_{1}k_{2}k}=0,k_{1},k_{2},k\mathrm{\ not\ all\ equal}\}$ must have $r_{kkk}=0$ for all $k$ (i.e. the first equation in Eqn. 4 cannot be satisfied).
Proof.
For any order-1 solution, for any $k$, in order to make $r_{k,-k,k}=z_{ak0}z_{b,-k,0}z_{ck0}=z_{ak0}\bar{z}_{bk0}z_{ck0}=0$, either $z_{ak0}$, $z_{bk0}$ or $z_{ck0}$ has to be zero, which means that $r_{kkk}=0$.
For order-2, first of all if any $z_{pk0}=0$ for any $p\in\{a,b,c\}$, then a constraint like $r_{k,k,-k}=z_{ak0}z_{bk0}\bar{z}_{ck0}+z_{ak1}z_{bk1}\bar{z}_{ck1}=0$ yields $z_{ak1}z_{bk1}z_{ck1}=0$ and thus $r_{kkk}=0$. If not, then for any two complex numbers $z_{pk0}$ and $z_{pk1}$, there always exist four real numbers $\theta_{p}\in(-\pi,\pi]$, $\theta^{\prime}_{p}\in(-\pi,\pi]$, $m_{p0}>0$ and $m_{p1}>0$ so that
$z_{pk0}=m_{p0}e^{\mathrm{i}\theta^{\prime}_{p}}e^{\mathrm{i}\theta_{p}},\quad\quad z_{pk1}=m_{p1}e^{\mathrm{i}\theta^{\prime}_{p}}e^{-\mathrm{i}\theta_{p}}$ (71)
Then a constraint like $r_{k,k,-k}=z_{ak0}z_{bk0}\bar{z}_{ck0}+z_{ak1}z_{bk1}\bar{z}_{ck1}=0$ can be written as $z_{ak0}z_{bk0}\bar{z}_{ck0}=-z_{ak1}z_{bk1}\bar{z}_{ck1}$, or equivalently:
$m_{a0}m_{b0}m_{c0}e^{\mathrm{i}(\theta^{\prime}_{a}+\theta^{\prime}_{b}+\theta^{\prime}_{c})}e^{\mathrm{i}(\theta_{a}+\theta_{b}-\theta_{c})}$ $=$ $-m_{a1}m_{b1}m_{c1}e^{\mathrm{i}(\theta^{\prime}_{a}+\theta^{\prime}_{b}+\theta^{\prime}_{c})}e^{-\mathrm{i}(\theta_{a}+\theta_{b}-\theta_{c})}$ (72)
$m_{a0}m_{b0}m_{c0}e^{\mathrm{i}\theta_{a}}e^{\mathrm{i}\theta_{b}}e^{-\mathrm{i}\theta_{c}}$ $=$ $-m_{a1}m_{b1}m_{c1}e^{-\mathrm{i}\theta_{a}}e^{-\mathrm{i}\theta_{b}}e^{\mathrm{i}\theta_{c}}$ (73)
Comparing their magnitude and phase, we have $m_{a0}m_{b0}m_{c0}=m_{a1}m_{b1}m_{c1}$ and
$\theta_{a}+\theta_{b}-\theta_{c}=\pm\pi/2\mod 2\pi$ (74)
Similarly, we have:
$\theta_{a}+\theta_{c}-\theta_{b}=\pm\pi/2\mod 2\pi,\quad\quad\theta_{b}+\theta_{c}-\theta_{a}=\pm\pi/2\mod 2\pi$ (75)
Solving the three equations and we have 6 possible solutions:
$(\theta_{a},\theta_{b},\theta_{c})$ $=(0,0,\pm\pi/2)\mod 2\pi$ (76)
$(\theta_{a},\theta_{b},\theta_{c})$ $=(0,\pm\pi/2,0)\mod 2\pi$ (77)
$(\theta_{a},\theta_{b},\theta_{c})$ =(±$=(\pm\pi/2,0,0)\mod 2\pi$ (78)
For all such solutions, let $m:=m_{a0}m_{b0}m_{c0}=m_{a1}m_{b1}m_{c1}$, then we have:
$r_{kkk}$ $=$ $z_{ak0}z_{bk0}z_{ck0}+z_{ak1}z_{bk1}z_{ck1}$ (79)
$=$ $me^{\mathrm{i}(\theta^{\prime}_{a}+\theta^{\prime}_{b}+\theta^{\prime}_{c})}(e^{\mathrm{i}(\theta_{a}+\theta_{b}+\theta_{c})}+e^{-\mathrm{i}(\theta_{a}+\theta_{b}+\theta_{c})})$ (80)
$=$ $me^{\mathrm{i}(\theta^{\prime}_{a}+\theta^{\prime}_{b}+\theta^{\prime}_{c})}(e^{\pm\mathrm{i}\pi/2}+e^{\mp\mathrm{i}\pi/2})$ (81)
$=$ 00 0 (82)
Lemma 6 (Property of order-3 solutions satisfying $R_{\mathrm{c}}$ and $R_{\mathrm{g}}$).
With small $L_{2}$ regularization, all per-frequency order-3 canonical solutions ${\bm{z}}$ at frequency $k_{0}$ that satisfy $R_{\mathrm{c}}$ and $R_{\mathrm{g}}$ are in the following form:
$z_{pk_{0}\cdot}=[1,\alpha_{p}\omega_{3},\beta_{p}\bar{\omega}_{3}],\quad\quad\mathrm{for\ }p\in\{a,b,c\}$ (83)
where $\alpha_{p}=\pm 1$ and $\beta_{p}=\pm 1$ with the constraint that $\alpha_{a}\alpha_{b}\alpha_{c}=\beta_{a}\beta_{b}\beta_{c}=1$. For $k\neq k_{0},z_{pk\cdot}=0$.
Proof.
We first prove that ${\bm{z}}$ satisfies $R_{\mathrm{c}}$ and $R_{\mathrm{g}}$. To see this, we have
$r_{k_{1}k_{2}k}$ $=\sum_{j}\mathbb{I}(k_{1}=k_{2}=k=k_{0})\omega_{3}^{3j}+\sum_{j}\mathbb{I}(-k_{1}=k_{2}=k=k_{0})\omega_{3}^{j}$ (84)
$+\ldots+\sum_{j}\mathbb{I}(-k_{1}=-k_{2}=-k=k_{0})\bar{\omega}_{3}^{3j}$ (85)
$=3\mathbb{I}(k_{1}=k_{2}=k=k_{0})+3\mathbb{I}(k_{1}=k_{2}=k=-k_{0})$ (86)
Note that all cross terms are gone since $\sum_{j}\omega^{j}_{3}=0$. It is clear that $r_{k_{1}k_{2}k}\neq 0$ unless $k_{1}=k_{2}=k$ so ${\bm{z}}$ satisfies $R_{\mathrm{c}}$ and $R_{\mathrm{g}}$.
Now we consider any per-frequency order-3 canonical solution (Def. 8) at frequency $k$. Let $a_{j}:=z_{akj}$, $b_{j}:=z_{bkj}$ and $c_{j}:=z_{ckj}$. Let ${\bm{a}}=[a_{j}]\in\mathbb{C}^{3}$, ${\bm{b}}=[b_{j}]\in\mathbb{C}^{3}$ and ${\bm{c}}=[c_{j}]\in\mathbb{C}^{3}$. Since the solution is canonical, we have $a_{0}=b_{0}=c_{0}=1$.
Then the conditions yield that
$({\bm{a}}\circ\bar{\bm{b}})^{\top}{\bm{c}}=0,\quad({\bm{a}}\circ\bar{\bm{b}})^{\top}\bar{\bm{c}}=0,\quad(\bar{\bm{a}}\circ{\bm{b}})^{\top}{\bm{c}}=0,\quad(\bar{\bm{a}}\circ{\bm{b}})^{\top}\bar{\bm{c}}=0$ (87)
which means that in $\mathbb{R}^{3}$ space, the following condition holds:
$\mathrm{span}(\Re({\bm{a}}\circ\bar{\bm{b}}),\Im({\bm{a}}\circ\bar{\bm{b}}))\perp\mathrm{span}(\Re({\bm{c}}),\Im({\bm{c}}))$ (88)
where $\Re(\cdot)$ and $\Im(\cdot)$ are real and imaginary parts of a complex vector. Since Eqn. 88 holds in $\mathbb{R}^{3}$, it must be the following cases: either $\Re({\bm{a}}\circ\bar{\bm{b}})$ is co-linear with $\Im({\bm{a}}\circ\bar{\bm{b}})$, or $\Re({\bm{c}})$ is co-linear with $\Im({\bm{c}})$.
If the latter is true (i.e., there exists $\beta$ so that $\beta\Re({\bm{c}})=\Im({\bm{c}})$), then since $c_{0}=1$ is real, $\beta=0$ and $\Im({\bm{c}})=0$. So ${\bm{c}}$ is real. In this case,
$r_{kkk}=({\bm{a}}\circ{\bm{b}})^{\top}{\bm{c}}=({\bm{a}}\circ{\bm{b}})^{\top}\bar{\bm{c}}=0$ (89)
If the former is true, then similarly we conclude that $\Im({\bm{a}}\circ\bar{\bm{b}})=0$ and ${\bm{a}}\circ\bar{\bm{b}}$ is real. Applying the same reasoning symmetrically, in order to find cases such that $r_{kkk}\neq 0$, a necessary condition is that
${\bm{a}}\circ\bar{\bm{b}},{\bm{b}}\circ\bar{\bm{c}},{\bm{c}}\circ\bar{\bm{a}}\in\mathbb{R}^{3}$ (90)
Let $z_{pkj}=|z_{pkj}|e^{\mathrm{i}\theta_{pj}}$. Let’s first consider the case that ${\bm{a}}\circ\bar{\bm{b}},{\bm{b}}\circ\bar{\bm{c}},{\bm{c}}\circ\bar{\bm{a}}\in\mathbb{R}^{3}_{\geq 0}$. Then we have $\theta_{a0}=\theta_{b0}=\theta_{c0}=\theta_{0}=0$, $\theta_{a1}=\theta_{b1}=\theta_{c1}=\theta_{1}$, $\theta_{a2}=\theta_{b2}=\theta_{c2}=\theta_{2}$. Letting $m_{j}:=|a_{j}||b_{j}||c_{j}|$, then the corresponding $r_{kkk}$ can be written as:
$r_{kkk}=\sum_{j=0}^{2}m_{j}e^{3\mathrm{i}\theta_{j}}$ (91)
with the constraints that $\sum_{j=0}^{2}m_{j}e^{\mathrm{i}\theta_{j}}=0$ imposed by $R_{\mathrm{c}}$.
Minimal Norm solutions. One interesting question is that what is the minimal norm representation that achieves the highest objective? For this we can solve the following optimization problem:
$\max_{\{m_{j},\theta_{j}\}}\sum_{j}m_{j}(e^{3\mathrm{i}\theta_{j}}+e^{-3\mathrm{i}\theta_{j}})-\epsilon\sum_{j}m^{2}_{j}\quad\mathrm{s.t.\ }\sum_{j}m_{j}e^{\mathrm{i}\theta_{j}}=0$ (92)
which achieves the maximal when $m_{j}=1/\epsilon$, $\theta_{1}=2\pi\mathrm{i}/3$ and $\theta_{2}=4\pi\mathrm{i}/3$ (or vise versa). Note that the optimal $\theta_{j}$ is fixed no matter how small the regularization coefficient $\epsilon$ is.
To see that, let $u_{j}:=e^{\mathrm{i}\theta_{j}}$. Then we have:
$\sum_{j}m_{j}(u_{j}+\bar{u}_{j})^{3}=\sum_{j}m_{j}[u^{3}_{j}+3u_{j}\bar{u}_{j}(u_{j}+\bar{u}_{j})+\bar{u}_{j}^{3}]=\sum_{j}m_{j}(u^{3}_{j}+\bar{u}^{3}_{j})$ (93)
Therefore, letting $x_{j}:=2\Re{u_{j}}$, we just need to consider the real part of the objective, and solve the following optimization in $\mathbb{R}$:
$\max_{\{m_{j},-2\leq x_{j}\leq 2,x_{0}=2\}}\sum_{j}m_{j}x^{3}_{j}-\epsilon\sum_{j}m^{2}_{j}\quad\mathrm{s.t.\ }\sum_{j}m_{j}x_{j}=0$ (94)
whose solutions give a sufficient condition. Using Lagrangian multiplier, we have:
$\frac{\partial L}{\partial x_{j}}=m_{j}(3x_{j}^{2}-\lambda)=0,\quad\quad\frac{\partial L}{\partial m_{j}}=x_{j}^{3}-2\epsilon m_{j}-\lambda x_{j}=0$ (95)
which leads to $\lambda=3$, $m_{j}=1/\epsilon$ and $x_{1}=x_{2}=-1$. This corresponds to the solution
$z_{pk\cdot}=[1,\omega_{3},\bar{\omega}_{3}],\quad\quad\mathrm{where\ }p\in\{a,b,c\}$ (96)
Note that the original necessary condition is ${\bm{a}}\circ\bar{\bm{b}},{\bm{b}}\circ\bar{\bm{c}},{\bm{c}}\circ\bar{\bm{a}}\in\mathbb{R}^{3}$. Considering the possible negativity, the solutions can be written as
$z_{pk\cdot}=[1,\alpha_{p}\omega_{3},\beta_{p}\bar{\omega}_{3}],\quad\quad\mathrm{for\ }p\in\{a,b,c\}$ (97)
where $\alpha_{p}=\pm 1$ and $\beta_{p}=\pm 1$ with the constraint that $\alpha_{a}\alpha_{b}\alpha_{c}=\beta_{a}\beta_{b}\beta_{c}=1$. ∎
Remarks. Note that this conclusion does not contradict with the constructed solution ${\bm{z}}_{\mathrm{syn},\alpha\beta}$ in Eqn. 11 in which $\alpha$ and $\beta$ are allowed to be any complex number with magnitude $1$. This is because ${\bm{z}}_{\mathrm{syn},\alpha\beta}$ does not satisfy all the constraints in $R_{\mathrm{c}}$ (but ${\bm{z}}_{\mathrm{syn},\alpha\beta}*{\bm{z}}_{\nu=\mathrm{i}}$ will) unless $\alpha$ and $\beta$ are real and thus ±$\pm 1$.